OpenStudy (goformit100):

2.75 + .003 + .158 = A. 4.36 B. 2.911 C. 0.436 D. 2.938 how to do it mentally ?

OpenStudy (goformit100):

@LoveYou*69

OpenStudy (anonymous):

2.911

OpenStudy (anonymous):

0.008+0.003+0.000 = 0.011 the answer must have 1 at the end.

OpenStudy (anonymous):

Keep them in the places you start them in: 2.75 0.003 0.158 8+3 = 11 so you carry the one over 5+5+1 = 11 so you carry it over again 7+1+1= 9 2+0+0 = 2 2.911

OpenStudy (goformit100):

Thanks